OpenStudy (anonymous):

Can someone help me factor s^3+5s^2+12s+8?

OpenStudy (anonymous):

Okay an easier question how do you get it to the form (s+1)(s^2+4s+8)

OpenStudy (zehanz):

It can be factored as (s - a)(s - b)(s - c), so s=a or s=b or s=c will make it zero. If you know one zero of s^3+5s^2+12s+8, you can find the others by dividing. To find a simple zero, just try some numbers. YOu'll see that s=-2 is a zero, so:\[s^3+5s^2+12s+8=(s+2)(2nd-degree-polynomial)\]is a partial factorisation. To find what 2nd degree polynomial it is, you can use synthetic division.

OpenStudy (anonymous):

I got two complex roots

OpenStudy (anonymous):

(s+1)(s+2+j2)(s+2-j2)

OpenStudy (anonymous):

Is there a method to getting it to (s+1)(s^2+4s+8)? without guessing to find a simple zero

OpenStudy (zehanz):

Sorry, typo! s = -1 is a zero, so you can factor out s+1. The 2nd degree polynomial turns out to be s²+4s+8. If you try to factor this further, you'll see the discriminant is negative, so there are no real zeros anymore...

OpenStudy (anonymous):

sorry how'd you get s=-1?

OpenStudy (zehanz):

See @allsmiles's answer. (Although I would use i instead of j)

OpenStudy (anonymous):

because it goes to 0 if you use -1?

OpenStudy (anonymous):

ahh I think I'm just slow but I still don't see how you ound out to use -1

OpenStudy (anonymous):

do you make the equation equal to 0 and then solve for s?

OpenStudy (zehanz):

yes, try it!

OpenStudy (anonymous):

I did I can't solve it you just get s^3+5s^2+12s+8=0

OpenStudy (anonymous):

I moved the 8 to the other side and factored the s out but I don't think thats the way to go about it

OpenStudy (zehanz):

What you have to realize is, in general you cannot easily factor a 3rd degree polynomial. If you are asked to do it, you can bet there are very "simple" numbers involved, like 1, -1, 2, -2 etc. In the case of s^3+5s^2+12s+8, trying s=-1 gives 0, so that is why you get (s+1)(2nd degree polynomial)

OpenStudy (anonymous):

alright I guess that will have to do for now thanks for your help

OpenStudy (zehanz):

Do you know synthetic division,? It sounds much more difficult than it is, actually! With it, you can get (s+1)(s²+4s+8) from s³+5s²+12s+8